OpenStudy (beefarm):

The electric field of an electromagnetic wave is given by; E=760sin(7.4x10^6pi(x-3x10^8t)) What is the wave's frequency? What is the amplitude of the magnetic field associated with this transverse electromagnetic wave?

OpenStudy (anonymous):

\[E=760 Sin(7.4\times10^{6}\pi x-2.22\times10^{15}\pi t)\] thus \[\omega=2\pi f = 2.22 \pi \times 10^{15} rad/s \rightarrow f=1.11\times10^{15}Hz\] \[B_{0}=E _{0}/c = 760V/3\times10^{8}ms ^{-1} = 2.5\times10^{-6} Tesla\]

OpenStudy (beefarm):

Thank you.

OpenStudy (anonymous):

My pleasure! I assumed that the E field is given in units of Volt/meter (V/m). If it's not, it will affect the units for B accordingly.